OpenStudy (spring98):

Part A: If (2^6)^x = 1, what is the value of x? Explain your answer. Part B: If (5^0)^x = 1, what are the possible values of x? Explain your answer.

OpenStudy (spring98):

@jdoe0001 Can you help me with this one too. I can't pas my quiz and if i don't get 80% or more i can't go on.

OpenStudy (jdoe0001):

\(\bf \textit{part A, multiply both sides by }\qquad \cfrac{1}{26} \\ \quad \\ \textit{part B, multiply both sides by }\qquad \cfrac{1}{50}\) what does that give you?

OpenStudy (spring98):

they both equal 1 @jdoe0001

OpenStudy (spring98):

is that right?

OpenStudy (jdoe0001):

\(\bf (26)x=1\implies \cfrac{1}{\cancel{26}}\cdot (\cancel{26})x=1\cdot \cfrac{1}{26}\implies x=? \\ \quad \\ \quad \\ (50)x=1\implies \cfrac{1}{\cancel{50}}\cdot (\cancel{50})x=1\cdot \cfrac{1}{50}\implies x=?\)

OpenStudy (spring98):

0.0384615384615385 is part A and 0.02for part B @jdoe0001

OpenStudy (jdoe0001):

\(\huge \circ.\circ\)

OpenStudy (jdoe0001):

notice, the number on the left cancelled out with the denominator

OpenStudy (spring98):

but what happened with the exponents?

OpenStudy (spring98):

@jdoe0001

OpenStudy (spring98):

@zepdrix @jdoe0001 can you guys plz help me?

OpenStudy (jdoe0001):

woops, got a typo there \(\bf (26)x=1\implies \cfrac{1}{\cancel{26}}\cdot (\cancel{26})x=1\cdot \cfrac{1}{26}\implies \cfrac{1}{1}\cdot x=\cfrac{1}{26}\implies x=\cfrac{1}{26} \\ \quad \\ \quad \\ (50)x=1\implies \cfrac{1}{\cancel{50}}\cdot (\cancel{50})x=1\cdot \cfrac{1}{50}\implies \cfrac{1}{1}\cdot x=\cfrac{1}{50}\implies x=\cfrac{1}{50}\)

OpenStudy (spring98):

i know but there is expopnents and i'm going to get this wrong.

OpenStudy (jdoe0001):

ohh...shoot.. is not what you originally posted though hmm.... need to dash in secs..... you may want to repost

zepdrix (zepdrix):

\[\large\rm (2^{6})^x=1\]Applying a rule of exponents to the left side gives us\[\large\rm 2^{6x}=1\]Rewrite the right side as a power of 2. 2 to what power = 1?

OpenStudy (spring98):

ok i will repost and i will tag you k @jdoe0001

OpenStudy (spring98):

Is it 0? @zepdrix

zepdrix (zepdrix):

Good. Anything to the zero power is 1. 2^0=1. Therefore we can write our equation:\[\large\rm 2^{6x}=1\]as\[\large\rm 2^{6x}=2^0\]Since the bases are the same, we can equate the exponents.

OpenStudy (spring98):

how did the 2 end up a 0. and isn't the x supposed to be a 0

OpenStudy (spring98):

@zepdrix

zepdrix (zepdrix):

I didn't change anything about the left side. I rewrite 1 as 2^0.

zepdrix (zepdrix):

Therefore, equating the exponents gives us \(\large\rm 6x=0\) which leads to \(\large\rm x=0\)

zepdrix (zepdrix):

For the other problem:\[\large\rm (5^0)^x = \color{orangered}{1}\]again, rewrite 1 as a power of 5,\[\large\rm (5^0)^x = \color{orangered}{5^0}\]Apply exponent rule to the left side,\[\large\rm (5^{0x}) = \color{orangered}{5^0}\]Which leads to\[\large\rm 0x=0\]And it looks like ANY value of x is going to solve this one.